How many grams of ammonium chloride must be added to 0.050 kg of water to produce a
solution that freezes at -5.00 °C? The cryoscopic constant of water is 1.86 °C kg mol-1 The
molar mass of ammonium chloride is 53.5 g mol-1.
1
Expert's answer
2020-06-10T07:38:54-0400

T2 = - 5.00 oC; T1 = 0.00 oC; Kf = 1.86 oC*kg/mol; mH2O = 0.050 kg; MNH4Cl = 53.5 g/mol

T1 - T2 = Kf*b

b = nNH4Cl/mH2O = mNH4Cl/(MNH4Cl*mH2O)

T1 - T2 = Kf*mNH4Cl/(MNH4Cl*mH2O)

mNH4Cl = MNH4Cl*mH2O*(T1 - T2)/Kf = 53.5*0.050*5/1.86 = 7.2 g